OpenStudy (anonymous):

(07.04) Fraction 4 over 5n = Fraction 2 over 3. n = ___? Fraction 2 over 15 Fraction 5 over 6 1Fraction 1 over 5 1Fraction 7 over 15

OpenStudy (igreen):

\(\dfrac{4}{5n} = \dfrac{2}{3}\) Cross multiply. 4 * 3 = 5n * 2

OpenStudy (anonymous):

its acualy \[\frac{ 4 }{ 5 }n\]

OpenStudy (anonymous):

No it isn't...

OpenStudy (anonymous):

Don't take advice from people who can't spell.

OpenStudy (anonymous):

yes it is

OpenStudy (igreen):

. . .

OpenStudy (anonymous):

Prove it.

OpenStudy (igreen):

Multiply 4 * 3

OpenStudy (anonymous):

it is just saying this is my q i should know what one of the numbers are

OpenStudy (anonymous):

12

OpenStudy (igreen):

Ohhh

OpenStudy (igreen):

Nvm

OpenStudy (anonymous):

Only listen to iGreen. He's the math champion.

OpenStudy (igreen):

So this is the question? \(\dfrac{4}{5}n = \dfrac{2}{3}\)

OpenStudy (anonymous):

and you do not know how to spell constantinople when you were in kinagarden so i do not wana her it

OpenStudy (igreen):

@OpiGeode Um..are you being sarcastic..?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Don't listen to the stuff pinkie pie says, it doesn't mean anything.

OpenStudy (anonymous):

and yes that is the question

OpenStudy (anonymous):

He/she is giving out wrong answers.

OpenStudy (anonymous):

i am standing here

OpenStudy (anonymous):

this is my question

OpenStudy (anonymous):

and i am a girl

OpenStudy (anonymous):

and that was not the answer that was one part of the question

OpenStudy (anonymous):

@OpiGeode

OpenStudy (anonymous):

@iGreen so can you write out the equation again correctly for me right now thats the only way i can funktion.

OpenStudy (igreen):

Okay, well we have to divide 4/5 to both sides, or multiply by 5/4. \(\dfrac{2}{3} \times \dfrac{5}{4} \rightarrow \dfrac{2 \times 5}{3 \times 4}\)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

20

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[\frac{ 10 }{ 12 }\]

OpenStudy (anonymous):

/2

OpenStudy (anonymous):

5 over 6

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

ok